OpenStudy (anonymous):

1. The length of Gordon Garden is 6 feet more than its width. A walkway 3 feet wide surrounds the outside of the garden. The total area of the garden and walkway is 720 square feet. Find the dimensions of the garden.

OpenStudy (cwrw238):

let x = width of garden so length = x + 6 total length including path = x + 6 + 2*3 = x + 12 .. width...................... = x + 2*3 = x + 6 total area = (x +6)(x+12) = 720 x^2 + 18x + 72 = 720 x^2 + 18x - 648 = 0 solve for x to get width and length

OpenStudy (cwrw238):

can you continue from here?

OpenStudy (anonymous):

not really

OpenStudy (cwrw238):

ok - this omes not straight forward you need 2 numbers whose product is -648 and whose sum is +18 you can find this from prime factors 18 = 2*3*3 648 =2*2*2*3*3*3*3 take out 2*3*3 out of factors of 648 leaves 2*2*3*3 = 36 - this gives 18 and 36 and +36 and -18 gives +18 - that gives factors of x^2 + 18x - 648: (x - 18)(x + 36) = 0 x = 18 or -36 and as - 36 cannot be the width width = 18 so length = 18 + 6 = 24 answer: dimensions of garden = 24 x 18